# Online Housing Market

### v1

@@ 6 Safe Serial Dictatorship Procedure
So far, I have considered and designed online algorithms consistent with most of the properties presented in Section 3, with the exception of $IR$. In this subsection, I explore the possibility to achieve $IR$ using the serial dictatorship procedure. Note that Proposition 4 states that the only algorithm achieving Pareto efficiency corresponds to Algorithm 1 (or equivalently Algorithm 2) using the ascending departure permutation $\delta$ as input. It is easy to check that this algorithm is neither $a$-$IC$, $d$-$IC$ nor $IR$, leading to the following corollary.
@@
This incompatibility with $IR$ is mainly due to the fact that in order to ensure Pareto efficiency in all cases, there is a need to risk at some point a violation of $IR$. In other words, among the full set of partial allocations, there are ones which may lead to a violation of $IR$ if no additional agents with appealing items appear later on. This risk can be highly beneficial if these hypothetical agents arrive with the right timing. However, this risk cannot be made in an online context to achieve $IR$ in all cases, since in the worst case, these agents will not come to counterbalance unfavorable exchanges. Therefore, in order to relax the notion of Pareto efficiency, I define a subset of allocations, called safe allocations, that will never lead to $IR$ violation in the following way.
@@
The definition of safe allocation means that after the departure of any agent, it is always possible to complete the current partial allocation to the agents that remain in the market, and to achieve individual rationality without considering agents that arrive later and their initial endowment. Algorithm 3 is an extension of Algorithm 2 that focuses on safe allocations instead of any possible allocations. By doing so, the partial allocation remains safe, and the algorithm is individually rational [^6] A similar algorithm has been considered in (?) where the agents choose sequentially their items, but the assignment problem is not online..
@@
Let me first show that Algorithm 3 using the ascending departure permutation $\delta$ as input is $IR$ and returns an $\mathcal{S}(I)$-$PO$ allocation for any instance $I$. The property of $IR$ is easy to show since, after each iteration of the “for” loop, the partial allocation $M$ can always be completed to be $IR$. This holds even if new agents arrive between two iterations, as their initial endowments can be assigned to them. Now, let me focus on the proof of the other property. The main argument is that the matching $M$ returned by the algorithm should be lexicographically optimal among the matchings in $\mathcal{S}(I)$, according to the order provided by the ascending departure permutation $\delta$ applied to $I$. Indeed, when the ascending departure permutation $\delta$ is used, there is no iteration during the while loop, and $T$ is empty when an agent chooses her item. Therefore, she chooses the best item among the items of $E_{<d_{\delta(i)}}\setminus A$ that maintains safeness. Let $\mathcal{P}$ denote the subset of safe allocations that may Pareto-dominate $M$. According to Algorithm 3, $M(\pi(1))$ should be the most favorite item of agent $\pi(1)$ in $E_{<d_{\pi(1)}}$ that maintains safeness. Therefore, all matchings in $\mathcal{P}\subseteq\mathcal{S}(I)$ should allocate $M(\pi(1))$ to agent $\pi(1)$ to Pareto-dominate $M$. Given that $M(\pi(1))$ should be allocated to agent $\pi(1)$, Algorithm 3 allocates to agent $\pi(2)$ her most favorite item in $E_{<d_{\pi(2)}}\setminus\{M(\pi(1))\}$ that maintains safeness. Therefore, all matchings in $\mathcal{P}$ should allocate $M(\pi(2))$ to agent $\pi(2)$ to Pareto-dominate $M$. By applying the same reasoning along the elements of permutation $\pi$, I can reach the conclusion that $\mathcal{P}$ may only contain a single allocation corresponding to $M$. However, a matching cannot Pareto-dominate itself, and therefore no matching of $\mathcal{S}(I)$ Pareto dominates $M$.
@@
Let me show now that any other online exchange algorithm cannot be both $IR$ and returns an $\mathcal{S}(I)$-$PO$ allocation for any instance $I$. I show this statement by contradiction by assuming that there exists another online exchange algorithm $\mathcal{A}$ that is both $IR$ and returns an $\mathcal{S}(I)$-$PO$ allocation for any instance $I$, and which behaves differently than Algorithm 3 using the ascending departure permutation $\delta$ as input. This means that there is an instance $I$ and an agent $j$ such that $\mathcal{A}_{j}(I)$ does not correspond to the assignment made by Algorithm 3 with the ascending departure permutation $\delta$ as input. In other words, agent $j$ does not receive her most favorite item among the remaining ones of $E_{<d_{\delta(j)}}$ that maintain safeness. Furthermore, I assume without loss of generality that agent $i$ is such an agent with the earliest departure time. Let $M$ denote the allocation obtained by Algorithm 3 applied to $I$ and using the ascending departure permutation $\delta$ as input. Note that if $\mathcal{A}_{j}(I)\succ_{j}M(j)$ holds, then $\mathcal{A}_{j}(I)$ is not safe and algorithm $\mathcal{A}$ is not $IR$ since $\mathcal{A}(I_{<d_{j}})$ cannot assign to each agent an item as preferred as her initial endowment, concluding the proof. Therefore, I can focus on the case where $M(j)\succ_{j}\mathcal{A}_{j}(I)$. Let $i$ denote the agent who receives $M(j)$ instead of agent $j$ in $\mathcal{A}(I)$, i.e., such that $\mathcal{A}_{i}(I)=M(j)$. By assumption, $d_{j}<d_{i}$ should hold, since otherwise an agent arriving earlier than agent $j$ would not receive the same item as in $M$.
@@
To show the contradiction, I construct a new instance $I^{\prime}$ by adding dummy agents to $I$ arriving later than the departure of agent $j$. For each agent $k$ leaving after $d_{j}$, let $k^{\prime}$ be a dummy agent such that $a_{k^{\prime}}=d_{j}+k\epsilon$, with $\epsilon>0$ small enough for each dummy agent to arrive earlier than any other departure, and $d_{k^{\prime}}$ large enough to leave after any other agent of $I$. The preferences of the agents are as follows. For any agent $k$ that was already in $I$ and leaving after $d_{j}$, her most favorite item becomes $e_{k^{\prime}}$, followed by the items of $I$ ranked in the same order as they were in her preferences in $I$, and finally all the other dummy items of $I^{\prime}$ ranked in arbitrary order. For dummy agent $i^{\prime}$ (corresponding to agent $i$), her preferences are $\mathcal{A}_{j}(I)\succ_{i^{\prime}}M(j)\succ_{i^{\prime}}\ldots$, where all the other items are ranked in arbitrary order. Finally, for any other dummy agent $k^{\prime}$, her most favorite item is $\mathcal{A}_{k}(I)$, and all other items are ranked arbitrarily.
@@
Note first that since $\mathcal{A}$ is online, both $\mathcal{A}_{j}(I^{\prime})=\mathcal{A}_{j}(I^{\prime}_{<d_{j}})$ and $\mathcal{A}_{j}(I_{<d_{j}})=\mathcal{A}_{j}(I)$ hold. Furthermore, since all dummy agents arrive later than $d_{j}$, $I_{<d_{j}}=I^{\prime}_{<d_{j}}$ holds, which implies with the former equalities that $\mathcal{A}_{j}(I^{\prime})=\mathcal{A}_{j}(I)$. Now concerning the assignment of the other agents, based on their preferences and since $\mathcal{A}$ returns an $\mathcal{S}(I^{\prime})$-$PO$ allocation, I can say that for each dummy agent $k^{\prime}$, I have $\mathcal{A}_{k^{\prime}}(I^{\prime})=\mathcal{A}_{k}(I)$, and for each regular agent $k$ leaving after $d_{j}$, I have $\mathcal{A}_{k}(I^{\prime})=e_{k^{\prime}}$. More precisely for $i^{\prime}$ I have $\mathcal{A}_{i^{\prime}}(I^{\prime})=\mathcal{A}_{i}(I)=M(j)$. Note that except for agents $i^{\prime}$ and $j$, any agent receives her most favorite item.
@@
But now consider the almost same allocation, say $M^{\prime}$, as the one constructed by $\mathcal{A}$, except that agent $j$ receives $M(j)$ and agent $i^{\prime}$ receives $\mathcal{A}_{j}(I)$. Note that by construction, $M^{\prime}$ should belong to $\mathcal{M}(I^{\prime})$ since $\mathcal{A}$ is an online exchange algorithm, and therefore $\mathcal{A}(I_{<d_{j}})=\mathcal{A}(I^{\prime}_{<d_{j}})$ belongs to $\mathcal{M}(I^{\prime}_{<d_{j}})$ (for agent $j$ and the ones leaving after her), $\mathcal{A}(I^{\prime})$ belongs to $\mathcal{M}(I^{\prime})$ (for all other agents except agent $i^{\prime}$), and $a_{i^{\prime}}<d_{i}$ holds (for agent $i^{\prime}$). Furthermore, $M^{\prime}$ belongs also to $\mathcal{S}(I^{\prime})$ since agent $j$ and each agent leaving before him receive the same item as in $\mathcal{A}(I)$, which is online, and during each later departure, the most favorite item of each agent remains in the market (corresponding to the ones they will ultimately receive). Agent $j$ and dummy agent $i^{\prime}$ receive in $M^{\prime}$ a strictly better item than the one offered by algorithm $\mathcal{A}$, since I have assumed that $M(j)\succ_{j}\mathcal{A}_{j}(I)$ holds. Therefore, $M^{\prime}$ Pareto dominates $\mathcal{A}(I^{\prime})$, leading to a contradiction since $\mathcal{A}$ should return an $\mathcal{S}(I^{\prime})$-$PO$ allocation. ∎
@@
Consider the instance described in the left part of Figure 3, as well as the allocations graphically described in the right part of the same figure. It is easy to check that the static serial dictatorship procedure with the ascending departure permutation $\delta$ produces allocation $M$, whereas the dynamic serial dictatorship procedure with the ascending arrival permutation $\alpha$ returns $M^{\prime}$.
Let me run the safe serial dictatorship with the ascending departure permutation $\delta$ on the same instance. Note first that this procedure, described in Algorithm 3, is essentially the same as Algorithm 2, with the exception that the set of items an agent can pick is restricted to those which leave the possibility for the remaining agents to be allocated items at least as favorable as their initial endowment. In other words, the set of items is restricted to those such that, if they are matched to the agent, it is still possible to complete the allocation to obtain a complete individually rational allocation for all agents present in the market. Therefore, at $d_{1}$, agent 1 is the first to choose, but she cannot pick item $e_{1}$ because otherwise it would be impossible to allocate the remaining item $e_{1}$ to agent 2 without violating individual rationality. Hence, item $e_{1}$ is allocated to agent 1. At $d_{3}$, agent 2 is the first to choose, and she reserves item $e_{3}$. Note that she can do so because agent 3 prefers the remaining item $e_{2}$ to her initial endowment $e_{3}$. Then, agent 3 picks $e_{2}$, which is allocated to her. At $d_{2}$, agent 2 picks item $e_{3}$ once again, which is allocated to her. Note that she can do so because item $e_{4}$ remains available for agent 4. Finally, at $d_{4}$, agent 4 picks the last remaining item $e_{4}$, which is allocated to her. The resulting allocation $M^{\prime\prime}$ is graphically described in Figure 3.
@@
Consider the following instance with three agents such that $a_{1}<a_{2}<a_{3}<d_{1}<d_{2}<d_{3}$. By contradiction, I assume that there is a permutation $\Pi$ that makes Algorithm 3 $WIC$. To simplify notation, I denote by $(i)$ the agent of rank $i$ in $\Pi(I)$ (instead of $\Pi_{i}(I)$). The preferences of the agents are provided depending on their rank in $\Pi(I)$ [^7] Note that I assume that the rank of an agent in a permutation function does not depend on her preferences., and are $e_{(3)}\succ_{(1)}e_{(2)}\succ_{(1)}e_{(1)}$, $e_{(1)}\succ_{(2)}e_{(3)}\succ_{(2)}e_{(2)}$, and $e_{(1)}\succ_{(3)}e_{(2)}\succ_{(3)}e_{(3)}$.
@@ 7 Top Trading Cycle Based Algorithms
In this section, I explore the possibility to extend the well-known top trading cycle algorithm to my online problem. The top trading cycle (TTC) procedure is a natural candidate to achieve the properties mentioned in Section 3, since in the offline context it fulfills all of them, and even stronger requirements (?, ?). Before providing an extension of TTC to the online setting, I provide a description of the standard TTC procedure in Algorithm 5, which will be used as a subroutine of my online algorithms. This procedure relies on the TTC-graph, which is an oriented graph where agents are vertices. As described in Algorithm 4, each vertex has an out-degree of one, and the single oriented edge starting at the vertex corresponding to an agent $i$ is oriented to the vertex corresponding to agent $j$ whose item $e_{j}$ is the most favorite according to $\succ_{i}$ among the ones detained by the agents in the graph. Such a graph should contain at least one cycle, and items are assigned along these cycles. Each agent contained in the cycle is assigned the item belonging to the agent that she is pointing to and leaves the market with this item. The procedure continues repeatedly with the remaining agents and their initial endowment until each agent is assigned an item.
@@
It is well known that the TTC procedure is $IR$ in the offline setting, and any online extension of this algorithm should keep this property. However, concerning the other properties, in the online setting not all agents are present at the beginning of the algorithm, and some of them will not arrive before the departure of other agents. Therefore, Algorithm 5 should be customized to fit the online setting. I explore a static (or greedy) procedure described in Algorithm 6, which applies TTC to subsets of agents already present in the market, according to an input partition function. As suggested by its name, partition function $\mathcal{P}$ applied to $I$ will partition the agents belonging to $I$ into multiple subsets. More formally, a partition function $\mathcal{P}:\mathcal{I}\to 2^{2^{N}}$ maps any instance $I\in\mathcal{I}$ to a partition $\mathcal{P}(I)$ of the agents in $I$. This means that $\mathcal{P}(I)$ is a collection of subsets of agents such that the union of these subsets is $N$, and no two subsets share an agent. Similarly to the permutation function used for the serial dictatorship procedure, I assume that the result of this function is independent of the preferences of the agents and their initial endowments. To construct the allocation, Algorithm 5 is applied independently to each subset of agents returned by this partition function. To simplify notations, I assume that $\mathcal{P}_{i}(I)$ denotes the subset of $\mathcal{P}(I)$ that contains agent $i$.
@@
A partition function $\mathcal{P}$ is static online compatible ($SOC$) if for any instance $I$ and agent $i$, $\mathcal{P}_{i}(I)=\mathcal{P}_{i}(I_{<d_{i}})$ holds.
Note that this property also implies that for any instance $I$ and for any agent $i$, $\mathcal{P}_{i}(I)$ contains only agents whose arrival times are no later than $d_{i}$. Furthermore, $\mathcal{P}_{i}(I)$ cannot contain an agent $j$ whose departure time $d_{j}$ is earlier than $a_{i}$, since otherwise I would have $\mathcal{P}_{i}(I_{<d_{i}})=\mathcal{P}_{i}(I)=\mathcal{P}_{j}(I)=\mathcal{P}_{j}(I_{<d_{j}})$, which is impossible, as agent $i$ is not part of $I_{<d_{j}}$. Therefore, $\mathcal{P}_{i}(I)$ contains only agents that are in the market at time $d_{i}$, i.e., for any $j\in\mathcal{P}_{i}(I)$, $a_{j}<d_{i}<d_{j}$ must hold. Following the same line of reasoning, it is easy to check that for any agent $j$ belonging to $\mathcal{P}_{i}(I_{<d_{i}})$ and leaving later than $d_{i}$, I have $\mathcal{P}_{j}(I_{<d_{j}})=\mathcal{P}_{j}(I_{<d_{i}})=\mathcal{P}_{i}(I_{<d_{i}})$. Therefore, no agent participates multiple times to the TTC procedure when the partition function used by Algorithm 6 is $SOC$. The following proposition shows that Algorithm 6 is online and agents have no incentive to misreport their preferences.
@@
Algorithm 6 using a $SOC$ partition function as input is both online and $WIC$.

### v2

@@ 6 Safe Serial Dictatorship Procedure
So far, we have considered and designed online algorithms consistent with most of the properties presented in Section 3, with the exception of $IR$. In this subsection, we explore the possibility to achieve $IR$ using the serial dictatorship procedure. Note that Proposition 4 states that the only algorithm achieving Pareto efficiency corresponds to Algorithm 1 (or equivalently Algorithm 2) using the ascending departure permutation $\delta$ as input. As suggested in Section 4.2 (see footnote 6), this algorithm is neither $a$-$IC$, $d$-$IC$ nor $IR$, leading to the following corollary.
@@
This incompatibility with $IR$ is mainly due to the fact that in order to ensure Pareto efficiency in all cases, there is a need to risk at some point a violation of $IR$. In other words, among the full set of partial allocations, there are ones which may lead to a violation of $IR$ if no additional agents with appealing items appear later on. This risk can be highly beneficial if these hypothetical agents arrive with the right timing. However, this risk cannot be made in an online context to achieve $IR$ in all cases, since in the worst case, these agents will not come to counterbalance unfavorable exchanges. Therefore, in order to relax the notion of Pareto efficiency, we define a subset of allocations, called safe allocations, that will never lead to $IR$ violation in the following way.
@@
The definition of safe allocation means that after the departure of any agent, it is always possible to complete the current partial allocation to the agents that remain in the market, and to achieve individual rationality without considering agents that arrive later and their initial endowment. Algorithm 3 is an extension of Algorithm 2 that focuses on safe allocations instead of any possible allocations. By doing so, the partial allocation remains safe, and the algorithm is individually rational. [^11] A similar algorithm was considered in (?), where agents select their items sequentially, but in an offline setting.
@@
Let us first show that Algorithm 3 using the ascending departure permutation $\delta$ as input is $IR$ and returns an $\mathcal{S}(I)$-$PO$ allocation for any instance $I$. The property of $IR$ is easy to show since, after each iteration of the “for” loop, the partial allocation $M$ can always be completed to be $IR$. This holds even if new agents arrive between two iterations, as their initial endowments can be assigned to them. Now, let us focus on the proof of the other property. The main argument is that the matching $M$ returned by the algorithm should be lexicographically optimal among the matchings in $\mathcal{S}(I)$, according to the order provided by the ascending departure permutation $\delta$ applied to $I$. Indeed, when the ascending departure permutation $\delta$ is used, there is no iteration during the while loop, and $T$ is empty when an agent chooses her item. Therefore, she chooses the best item among the items of $E_{<d_{\delta(i)}}\setminus A$ that maintains safeness. Let $\mathcal{P}$ denote the subset of safe allocations that may Pareto-dominate $M$. According to Algorithm 3, $M(\pi(1))$ should be the most favorite item of agent $\pi(1)$ in $E_{<d_{\pi(1)}}$ that maintains safeness. Therefore, all matchings in $\mathcal{P}\subseteq\mathcal{S}(I)$ should allocate $M(\pi(1))$ to agent $\pi(1)$ to Pareto-dominate $M$. Given that $M(\pi(1))$ should be allocated to agent $\pi(1)$, Algorithm 3 allocates to agent $\pi(2)$ her most favorite item in $E_{<d_{\pi(2)}}\setminus\{M(\pi(1))\}$ that maintains safeness. Therefore, all matchings in $\mathcal{P}$ should allocate $M(\pi(2))$ to agent $\pi(2)$ to Pareto-dominate $M$. By applying the same reasoning along the elements of permutation $\pi$, we can reach the conclusion that $\mathcal{P}$ may only contain a single allocation corresponding to $M$. However, a matching cannot Pareto-dominate itself, and therefore no matching of $\mathcal{S}(I)$ Pareto dominates $M$.
@@
Let us show now that any other online exchange algorithm cannot be both $IR$ and returns an $\mathcal{S}(I)$-$PO$ allocation for any instance $I$. We show this statement by contradiction by assuming that there exists another online exchange algorithm $\mathcal{A}$ that is both $IR$ and returns an $\mathcal{S}(I)$-$PO$ allocation for any instance $I$, and which behaves differently than Algorithm 3 using the ascending departure permutation $\delta$ as input. This means that there is an instance $I$ and an agent $j$ such that $\mathcal{A}_{j}(I)$ does not correspond to the assignment made by Algorithm 3 with the ascending departure permutation $\delta$ as input. In other words, agent $j$ does not receive her most favorite item among the remaining ones of $E_{<d_{\delta(j)}}$ that maintain safeness. Furthermore, we assume without loss of generality that agent $i$ is such an agent with the earliest departure time. Let $M$ denote the allocation obtained by Algorithm 3 applied to $I$ and using the ascending departure permutation $\delta$ as input. Note that if $\mathcal{A}_{j}(I)\succ_{j}M(j)$ holds, then $\mathcal{A}_{j}(I)$ is not safe and algorithm $\mathcal{A}$ is not $IR$ since $\mathcal{A}(I_{<d_{j}})$ cannot assign to each agent an item as preferred as her initial endowment, concluding the proof. Therefore, we can focus on the case where $M(j)\succ_{j}\mathcal{A}_{j}(I)$. Let $i$ denote the agent who receives $M(j)$ instead of agent $j$ in $\mathcal{A}(I)$, i.e., such that $\mathcal{A}_{i}(I)=M(j)$. By assumption, $d_{j}<d_{i}$ should hold, since otherwise an agent arriving earlier than agent $j$ would not receive the same item as in $M$.
@@
To show the contradiction, we construct a new instance $I^{\prime}$ by adding dummy agents to $I$ arriving later than the departure of agent $j$. For each agent $k$ leaving after $d_{j}$, let $k^{\prime}$ be a dummy agent such that $a_{k^{\prime}}=d_{j}+k\epsilon$, with $\epsilon>0$ small enough for each dummy agent to arrive earlier than any other departure, and $d_{k^{\prime}}$ large enough to leave after any other agent of $I$. The preferences of the agents are as follows. For any agent $k$ that was already in $I$ and leaving after $d_{j}$, her most favorite item becomes $e_{k^{\prime}}$, followed by the items of $I$ ranked in the same order as they were in her preferences in $I$, and finally all the other dummy items of $I^{\prime}$ ranked in arbitrary order. For dummy agent $i^{\prime}$ (corresponding to agent $i$), her preferences are $\mathcal{A}_{j}(I)\succ_{i^{\prime}}M(j)\succ_{i^{\prime}}\ldots$, where all the other items are ranked in arbitrary order. Finally, for any other dummy agent $k^{\prime}$, her most favorite item is $\mathcal{A}_{k}(I)$, and all other items are ranked arbitrarily.
@@
Note first that since $\mathcal{A}$ is online, both $\mathcal{A}_{j}(I^{\prime})=\mathcal{A}_{j}(I^{\prime}_{<d_{j}})$ and $\mathcal{A}_{j}(I_{<d_{j}})=\mathcal{A}_{j}(I)$ hold. Furthermore, since all dummy agents arrive later than $d_{j}$, $I_{<d_{j}}=I^{\prime}_{<d_{j}}$ holds, which implies with the former equalities that $\mathcal{A}_{j}(I^{\prime})=\mathcal{A}_{j}(I)$. Now concerning the assignment of the other agents, based on their preferences and since $\mathcal{A}$ returns an $\mathcal{S}(I^{\prime})$-$PO$ allocation, we can say that for each dummy agent $k^{\prime}$, we have $\mathcal{A}_{k^{\prime}}(I^{\prime})=\mathcal{A}_{k}(I)$, and for each regular agent $k$ leaving after $d_{j}$, we have $\mathcal{A}_{k}(I^{\prime})=e_{k^{\prime}}$. More precisely for $i^{\prime}$ we have $\mathcal{A}_{i^{\prime}}(I^{\prime})=\mathcal{A}_{i}(I)=M(j)$. Note that except for agents $i^{\prime}$ and $j$, any agent receives her most favorite item.
@@
But now consider the almost same allocation, say $M^{\prime}$, as the one constructed by $\mathcal{A}$, except that agent $j$ receives $M(j)$ and agent $i^{\prime}$ receives $\mathcal{A}_{j}(I)$. Note that by construction, $M^{\prime}$ should belong to $\mathcal{M}(I^{\prime})$ since $\mathcal{A}$ is an online exchange algorithm, and therefore $\mathcal{A}(I_{<d_{j}})=\mathcal{A}(I^{\prime}_{<d_{j}})$ belongs to $\mathcal{M}(I^{\prime}_{<d_{j}})$ (for agent $j$ and the ones leaving after her), $\mathcal{A}(I^{\prime})$ belongs to $\mathcal{M}(I^{\prime})$ (for all other agents except agent $i^{\prime}$), and $a_{i^{\prime}}<d_{i}$ holds (for agent $i^{\prime}$). Furthermore, $M^{\prime}$ belongs also to $\mathcal{S}(I^{\prime})$ since agent $j$ and each agent leaving before him receive the same item as in $\mathcal{A}(I)$, which is online, and during each later departure, the most favorite item of each agent remains in the market (corresponding to the ones they will ultimately receive). Agent $j$ and dummy agent $i^{\prime}$ receive in $M^{\prime}$ a strictly better item than the one offered by algorithm $\mathcal{A}$, since we have assumed that $M(j)\succ_{j}\mathcal{A}_{j}(I)$ holds. Therefore, $M^{\prime}$ Pareto dominates $\mathcal{A}(I^{\prime})$, leading to a contradiction since $\mathcal{A}$ should return an $\mathcal{S}(I^{\prime})$-$PO$ allocation. ∎
@@
Consider the instance described in the left part of Figure 3, as well as the allocations graphically described in the right part of the same figure. It is easy to check that the static serial dictatorship procedure (Algorithm 1) with the ascending departure permutation $\delta$ produces allocation $M$, whereas the dynamic serial dictatorship procedure (Algorithm 2) with the ascending arrival permutation $\alpha$ returns $M^{\prime}$.
Let us run the safe serial dictatorship with the ascending departure permutation $\delta$ on the same instance. Note first that this procedure, described in Algorithm 3, is essentially the same as Algorithm 2, with the exception that the set of items an agent can pick is restricted to those which leave the possibility for the remaining agents to be allocated items at least as favorable as their initial endowment. In other words, the set of items is restricted to those such that, if they are matched to the agent, it is still possible to complete the allocation to obtain a complete individually rational allocation for all agents present in the market. Therefore, at $d_{1}$, agent 1 is the first to choose, but she cannot pick item $e_{2}$ because otherwise it would be impossible to allocate the remaining item $e_{1}$ to agent 2 without violating individual rationality. Hence, item $e_{1}$ is allocated to agent 1. At $d_{3}$, agent 2 is the first to choose, and she reserves item $e_{3}$. Note that she can do so because agent 3 prefers the remaining item $e_{2}$ to her initial endowment $e_{3}$. Then, agent 3 picks $e_{2}$, which is allocated to her. At $d_{2}$, agent 2 picks item $e_{3}$ once again, which is allocated to her. Note that she can do so because item $e_{4}$ remains available for agent 4. Finally, at $d_{4}$, agent 4 picks the last remaining item $e_{4}$, which is allocated to her. The resulting allocation $M^{\prime\prime}$ is graphically described in Figure 3.
@@
Consider the following instance with three agents such that $a_{1}<a_{2}<a_{3}<d_{1}<d_{2}<d_{3}$. By contradiction, we assume that there is a permutation $\Pi$ that makes Algorithm 3 $WIC$. To simplify notation, we denote by $(i)$ the agent of rank $i$ in $\Pi(I)$ (instead of $\Pi_{i}(I)$). The preferences of the agents are provided depending on their rank in $\Pi(I)$ [^12] Note that we assume that the rank of an agent in a permutation function does not depend on her preferences., and are $e_{(3)}\succ_{(1)}e_{(2)}\succ_{(1)}e_{(1)}$, $e_{(1)}\succ_{(2)}e_{(3)}\succ_{(2)}e_{(2)}$, and $e_{(1)}\succ_{(3)}e_{(2)}\succ_{(3)}e_{(3)}$.
@@ 7 Top Trading Cycle Based Algorithms
In this section, we explore the possibility to extend the well-known top trading cycle algorithm to our online problem (?). The top trading cycle (TTC) procedure is a natural candidate to achieve the properties mentioned in Section 3, since in the offline context it fulfills all of them, and even stronger requirements (?, ?). Before providing an extension of TTC to the online setting, we provide a description of the standard TTC procedure in Algorithm 5, which will be used as a subroutine of our online algorithms. This procedure relies on the TTC-graph, which is an oriented graph where agents are vertices. As described in Algorithm 4, which explains its construction, each vertex has out-degree one: the unique outgoing edge from the vertex corresponding to agent $i$ points to the vertex corresponding to the agent $j$ whose item $e_{j}$ is, according to $\succ_{i}$, the most preferred among those held by the agents currently in the graph. Such a graph should contain at least one cycle, and items are assigned along these cycles. Each agent contained in the cycle is assigned the item belonging to the agent that she is pointing to and leaves the market with this item. The procedure continues repeatedly with the remaining agents and their initial endowment until each agent is assigned an item.
@@
It is well known that the TTC procedure is $IR$ in the offline setting, and any online extension of this algorithm should keep this property. However, concerning the other properties, in the online setting not all agents are present at the beginning of the algorithm, and some of them will not arrive before the departure of other agents. Therefore, Algorithm 5 should be customized to fit the online setting. We explore a static (or greedy) procedure described in Algorithm 6, which applies TTC to subsets of agents already present in the market, according to an input partition function. As suggested by its name, partition function $\mathcal{P}$ applied to $I$ will partition the agents belonging to $I$ into multiple subsets. More formally, a partition function $\mathcal{P}:\mathcal{I}\to 2^{2^{N}}$ maps any instance $I\in\mathcal{I}$ to a partition $\mathcal{P}(I)$ of the agents in $I$. This means that $\mathcal{P}(I)$ is a collection of subsets of agents such that the union of these subsets is $N$, and no two subsets share an agent. Similarly to the permutation function used in the serial dictatorship procedure, we assume that its output is independent of the agents’ preferences and initial endowments, and depends only on their arrival and departure times. To construct the allocation, Algorithm 5 is applied independently to each subset of agents returned by this partition function. To simplify notations, we assume that $\mathcal{P}_{i}(I)$ denotes the subset of $\mathcal{P}(I)$ that contains agent $i$.
@@
A partition function $\mathcal{P}$ is progress-preserving ($PP$) if for any instance $I$ and agent $i$, $\mathcal{P}_{i}(I)=\mathcal{P}_{i}(I_{<d_{i}})$ holds.
Note that this property also implies that for any instance $I$ and for any agent $i$, $\mathcal{P}_{i}(I)$ contains only agents whose arrival times are no later than $d_{i}$. Furthermore, $\mathcal{P}_{i}(I)$ cannot contain an agent $j$ whose departure time $d_{j}$ is earlier than $a_{i}$, since otherwise we would have $\mathcal{P}_{i}(I_{<d_{i}})=\mathcal{P}_{i}(I)=\mathcal{P}_{j}(I)=\mathcal{P}_{j}(I_{<d_{j}})$, which is impossible, as agent $i$ is not part of $I_{<d_{j}}$. Therefore, $\mathcal{P}_{i}(I)$ contains only agents that are in the market at time $d_{k}=\min_{j\in\mathcal{P}_{i}(I)}d_{j}$, i.e., for any $j\in\mathcal{P}_{i}(I)$, $a_{j}<d_{k}\leq d_{j}$ must hold. Following the same line of reasoning, it is easy to check that for any agent $j$ belonging to $\mathcal{P}_{i}(I_{<d_{i}})$ and leaving later than $d_{i}$, we have $\mathcal{P}_{j}(I_{<d_{i}})=\mathcal{P}_{i}(I_{<d_{i}})=\mathcal{P}_{i}(I_{<d_{j}})$, which implies that $\mathcal{P}_{j}(I_{<d_{j}})=\mathcal{P}_{i}(I_{<d_{j}})$ since $j\in\mathcal{P}_{j}(I_{<d_{i}})=\mathcal{P}_{i}(I_{<d_{j}})$. Therefore, no agent participates multiple times to the TTC procedure when the partition function used by Algorithm 6 is $PP$. The following proposition shows that Algorithm 6 is online and agents have no incentive to misreport their preferences.
@@
Algorithm 6 using a $PP$ partition function as input is both online and $WIC$.